Weiland Co. shows the following information on its 2016 income statement sales $160,500 costs $80.400, other expenses $3,700; depreciation expense $9,400 interest expense $6,900, taxes $21,035, dividends $7,950. In addition, youre told that the frm issued $4,100 in new ecuity during 2016 and redeemed $6,900 in outstanding long-term debt a. What is the 2016 operating cash flow? (Do not round intermediate calculations and round your answer to the nearest whole sumber, e.g. 32.) Operating cash flow b. What is the 2016 cash flow to creditors? (Do not round intermediate calculations and round your answer to the nearest whole number,g,32) Cash flow to creditors e. What is the 2016 cash fow to stockholders? (Do not round intermediate calculations and round your answer to the nearest whole number,e.g 32) Cash flow to stockholders d. If net fxed assets increased by $21,000 during the year, what was the addtion to NWC? (Do not round intermediate calculations and round your answer to the nearest whole number,.932.) Addtion to net working capital $
Answer #1
Answer a)
operating cash flow = Income after tax +Interest + depreciation
Income after tax =
Sales 160500
Less Cost 80400
Less Other expenses 3700
Less Depreciation expenses 9400
Less Interest expenses 6900
Less tax 21035
Income after tax 39065
Operating cash flow
Income after tax 39065
Add Depreciation expenses 9400
Add Interest expenses 6900
Operating cash flow 55365
Answer b)
Cash flow to creditor = Interest paid + Debt redeeption
Cash flow to creditor = =6900+6900
Cash flow to creditor = 13800
Answer c)
Cash flow to stockholder
Cash flow to stockholder = Dividend paid - new equity issue
Cash flow to stockholder =7950-4100
Cash flow to stockholder 3850
Answer d)
Change in NWC

Net capital spending = depreciation+increase in FA
Net capital spend = 9400+21000
Net capital spend = 30400
Cash flow from assets=cash flow to creditors+cash flow to stockholders
Cash flow from asset =13800+3850
Cash flow from asset (CFFA) 17650
CFFA=OCF-Change in NWC-net capital spending
=17650=55365-Change in NWC-30400
Change in NWC = =55365-30400-17650
Change in NWC = 7315